According to BLS news release, in December 2019 the unemployment rate and the labor force participation rate were 3.5% and 63.2 respectively. What would be the unemployment rate if the participation rate were 66%?
Answer #1

Let the population be x

In case participation rate is 63.2%,

Labor force=LF=x*63.2%=0.632x

Number of unemployed persons=LF*Unemployment rate=0.632x*3.5%=0.02212x

In case participation rate is 66%

Labor force=LF=x*66%=0.66x

Unemployment rate=Number of unemployed persons/LF=0.02212x/0.66x=3.35%